# On a classification of Morse functions on $3$-dimensional manifolds represented as connected sums of manifolds of Heegaard genus one

### v1

@@ Abstract.
    On the edge set $E_{K}$ of $K$, a map $l_{K}:E_{K}\rightarrow\{0,1\}$ is given and the size of the set ${l_{K}}^{-1}(1)$ is $c$.
@@
Conversely, for such a function on a $3$-dimensional closed and orientable manifold $M$ with such a Reeb digraph $W_{f}$, $M$ is diffeomorphic to one of these manifolds. Here the case $r=0$ and $c^{\prime}=0$ implies $M$ is diffeomorphic to $S^{3}$.
@@
Let $d$ be the number of handles and singular points here. We choose an order on the set of the $d$ singular points of the function. We can also deform the function into another Morse function by a suitable homotopy without changing its singular points and the values on $c^{-1}(\mathbb{R}-(q-\epsilon,q+\epsilon))$ in such a way that at distinct singular points of the resulting new function $c^{\prime}:X\rightarrow\mathbb{R}$ the values are distinct. More precisely, we can change the function in such a way that at the $j$-th singular point of the $d$ singular points ($1\leq j\leq d$) of the function, the value is $q_{j}$ with $q-\epsilon<q_{j}<q+\epsilon$ and that the relation $q_{j_{1}}<q_{j_{2}}$ is satisfied for $1\leq j_{1}<j_{2}\leq d$. We can also change the order on the set $\{s_{1},\cdots,s_{j}\}$ in arbitrary ways (there exist $d!$ cases). For each singular point of the new function, the handle same as that in the original function corresponds to. In other words, we can attach handles one after another to the (original) mutually disjoint submanifolds diffeomorphic to $\partial D^{k}\times D^{m-k-1}$ of the preimage of $q-\epsilon$, according to the values, increasing from $q-\epsilon$ to $q+\epsilon$.
@@
We can know this as a kind of elementary exercise on Morse functions, singular points of them, and handles. Hereafter we call such a Morse function a simple sphere-torus-fibered Morse function or an SSTF Morse function.
@@
This reconstruction is first presented in [15] in the case preimages of single points containing no singular points are disjoint unions of spheres and later in [7] for our case. As more generalized cases, our studies [9, 10, 11] for Morse functions on manifolds whose dimensions are greater than $2$ and preimages of single points are also respected: however we do not assume related knowledge and arguments at all.
@@
    For a Morse function explained just before, this can be corresponded to some attachment of handles to the sphere ${f_{0,v}}^{-1}(t_{v,1}+\epsilon)$ to obtain $X_{v}={f_{0,v}}^{-1}([t_{v,1},t_{v,2}])$ according to the following procedure where $\epsilon>0$ is a sufficiently small number. This respects [saek1, Lemma 6.6], presented as Theorem 5 later, and its original proof. More precisely, in Theorem 5, we consider the case $g=0$. We argue our handle attachment.
@@
      We attach each $1$-handle to a disjoint union of two $2$-dimensional disks of these all $2k_{2}$ copies of $2$-dimensional disks to connect the $k_{1}+1$ copies of the sphere $S^{2}\subset S_{v,1}$. We can connect the disjoint union $S_{v,1}$ of spheres to a sphere only in this way. In addition, the relation $k_{1}=k_{2}:=k$ must hold.
@@
    At each remaining vertex $v$, we can locally have a Morse function $f_{v}:X_{v}\rightarrow\mathbb{R}$ on a $3$-dimensional compact and orientable manifold $X_{v}$ enjoying the following properties as an example. Let $i_{v,S^{2}}\geq 0$ ($i_{v,S^{1}\times S^{1}}\geq 0$) denote the number of oriented edges entering $v$ where $l_{K}$ has the value 00 (resp. $1$). Let $o_{v,S^{2}}\geq 0$ ($o_{v,S^{1}\times S^{1}}\geq 0$) denote the number of oriented edges departing from $v$ where $l_{K}$ has the value 00 (resp. $1$). The former case is for $i_{v,S^{2}}=o_{v,S^{2}}=1$ and $i_{v,S^{1}\times S^{1}}=o_{v,S^{1}\times S^{1}}=0$ and we do not consider this case here.
@@
  • First we can see that for our resulting SSTF Morse function $f_{0}$ there exists exactly $c$ edges of $W_{f_{0}}$ such that the preimage ${q_{f_{0}}}^{-1}(t)$ of $t$ in the interior is diffeomorphic to the torus $S^{1}\times S^{1}$. Such edges are not mutually adjacent by our construction. The 1st Betti number of $W_{f_{0}}$ is greater than or equal to $a+b$. We can make the 1st Betti number an arbitrary integer $r\geq a+b$ by our construction. We remove the preimage of a small connected regular neighborhood of each of these $c$ edges. We have a $3$-dimensional manifold whose boundary is diffeomorphic to a disjoint union of $2c$ copies of $S^{2}$. We attach copies of the cylinder $S^{2}\times D^{1}$ instead to the boundary. This naturally yields another SSTF Morse function such that preimages of single points containing no singular points of it are disjoint unions of spheres. The resulting $3$-dimensional closed and connected manifold is diffeomorphic to the sphere $S^{3}$ or a manifold diffeomorphic to a connected sum $\sharp(S^{1}\times S^{2})$ of $r\geq a+b$ copies of $S^{1}\times S^{2}$. This is regarded as a higher dimensional version of Theorem 1. We can also understand the topology of the manifold $M$ from classical $3$-dimensional manifold theory. We may also have various ways of understanding. For example, on fundamental groups, we can also apply isomorphisms of fundamental groups of the manifold and the Reeb graph, discussed in [22, Corollary 4.8], [7], and [8, Corollary 4].
@@
We can deform such a Morse function to an SSTF Morse function $f_{0}:M\rightarrow\mathbb{R}$ by a suitable homotopy. We can also have such a function $f_{0}$ in such a way that to each local SSTF Morse function $f_{0,v}:X_{v}\rightarrow\mathbb{R}$ around each vertex $v$, the following handle attachment can be corresponded. Let $f(X_{v})=[t_{v,1},t_{v,2}]$ as in STEP 2.
@@
This is thanks to the topology of the manifold and fundamental arguments on $3$-dimensional manifold theory. The $3$-dimensional manifolds of the preimages of the local graphs in FIGURE 5 are all diffeomorphic to $(S^{1}\times D^{2})-{\rm Int}\ (D^{2}\sqcup D^{2})$. Here ${\rm Int}\ X$ means the interior $X-\partial X$ of the topological manifold $X$, the two disjoint disks $D^{2}\sqcup D^{2}$ are smoothly embedded in the interior of $S^{1}\times D^{2}$ and here its interior is removed. Note that similar local changes are presented in [16, Figure 4] and also used in [13] and that we cannot show this argument on the $3$-dimensional manifold from them directly. More precisely, we can realize the local movements by applying suitable operations in [16, Figure 5] twice. This gives another proof for the local movements.
@@
Let a conected digraph $K_{g}$ be defined from a piecewise smooth function $g:K\rightarrow\mathbb{R}$ with the condition that the degree of a vertex where the function $g$ has a local extremum is of degree at most $2$. Then we have the following.
@@
    At distinct singular points of $f_{\rm m}$ the values are distinct.
@@
The author would like to thank members of the research group of Osamu Saeki and people organizing and supporting Saga Souhatsu Mathematical Seminar (http://inasa.ms.saga-u.ac.jp/Japanese/saga-souhatsu.html): the author would also like to thank Inasa Nakamura for inviting the author as a speaker and letting him present [9, 11]. Discussing several Morse functions and topological properties of the manifolds of their domains with them has motivated the author to challenge the present study.
@@
The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/) and this study is closely related to our study. Our study thanks them for the supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. Although he is not employed there, our study also thanks this.
Some work by other researchers and this version may overlap in some of the contents due to the nature that our problems are natural in theory of Morse functions and applications to differential topology and that related mathematical studies are very fundamental and classical in some senses, for example. However the present version of our paper is presented independent of these work.

### v2

@@ Abstract.
    A map $l_{K}:E_{K}\rightarrow\{0,1\}$ on the edge set $E_{K}$ of $K$ satisfies that the number of vertices $v$ of degree $2$ contained in two edges $e_{v,1}$ and $e_{v,2}$ with $l_{K}(e_{v,1})=l_{K}(e_{v,2})=0$ is $b$ and that the size of the set ${l_{K}}^{-1}(1)$ is $c$.
@@
Conversely, for a Morse function $f:M\rightarrow\mathbb{R}$ on a $3$-dimensional closed and orientable manifold $M$ with such a Reeb digraph $W_{f}$, $M$ is diffeomorphic to one of these manifolds. Here the case $r=0$ and $c^{\prime}=0$ implies $M$ is diffeomorphic to $S^{3}$.
@@
Let $d$ be the number of handles and singular points of the function here. We choose an order on the set of the $d$ singular points of the function. We can also deform the function into another Morse function by a suitable homotopy without changing its singular points and the values on $c^{-1}(\mathbb{R}-(q-\epsilon,q+\epsilon))$ in such a way that at distinct singular points of the resulting new function $c^{\prime}:X\rightarrow\mathbb{R}$ the values are distinct. More precisely, we can change the function in such a way that at the $j$-th singular point of the $d$ singular points ($1\leq j\leq d$) of the function, the value is $q_{j}$ with $q-\epsilon<q_{j}<q+\epsilon$ and that the relation $q_{j_{1}}<q_{j_{2}}$ is satisfied for $1\leq j_{1}<j_{2}\leq d$. We can also change the order on the set $\{s_{1},\cdots,s_{j}\}$ in arbitrary ways (there exist $d!$ cases). For each singular point of the new function, the handle same as that in the original function corresponds to. In other words, we can attach handles one after another to the (original) mutually disjoint submanifolds diffeomorphic to $\partial D^{k}\times D^{m-k-1}$ of the preimage of $q-\epsilon$, according to the values, increasing from $q-\epsilon$ to $q+\epsilon$.
@@
We can know this as a kind of elementary exercise on Morse functions, singular points of them, and handles.
Hereafter we call such a Morse function a simple sphere-torus-fibered Morse function or an SSTF Morse function.
Hereafter we call a Morse function such that at distinct singular points of the function the values are always distinct a simple Morse function. An SSTF Morse function is simple of course.
@@
This reconstruction is first presented in [15] in the case preimages of single points containing no singular points of the functions are disjoint unions of spheres. Later, in [7], the case preimages of single points containing no singular points of the functions are disjoint unions of (2-dimensional) spheres and closed, connected and orientable surfaces, is studied. As more generalized cases, we also respect our studies [9, 10, 11] for Morse functions on manifolds whose dimensions are greater than $2$ and preimages of single points diffeomorphic to more general manifolds: however we do not assume related knowledge and arguments at all.
@@
    For a Morse function explained just before, this can be corresponded to some attachment of handles to the sphere ${f_{0,v}}^{-1}(t_{v,1}+\epsilon)$ to obtain $X_{v}={f_{0,v}}^{-1}([t_{v,1},t_{v,2}])$ according to the following procedure where $\epsilon>0$ is a sufficiently small number. This respects [20, Lemma 6.6], presented as Theorem 5 later, and its original proof. More precisely, in Theorem 5, we consider the case $g=0$. We argue our handle attachment.
@@
      We attach each $1$-handle to a disjoint union of two $2$-dimensional disks of these all $2k_{2}$ copies of $2$-dimensional disks to connect the $k_{1}+1$ copies of the sphere $S^{2}\subset S_{v,1}$. We can connect the disjoint union $S_{v,1}$ of spheres to obtain a sphere as a result only in this way. In addition, the relation $k_{1}=k_{2}:=k$ must hold.
@@
    At each remaining vertex $v$, we can locally have a Morse function $f_{v}:X_{v}\rightarrow\mathbb{R}$ on a $3$-dimensional compact and orientable manifold $X_{v}$ enjoying the following properties as an example. Let $i_{v,S^{2}}\geq 0$ ($i_{v,S^{1}\times S^{1}}\geq 0$) denote the number of oriented edges entering $v$ where $l_{K}$ has the value 00 (resp. $1$). Let $o_{v,S^{2}}\geq 0$ ($o_{v,S^{1}\times S^{1}}\geq 0$) denote the number of oriented edges departing from $v$ where $l_{K}$ has the value 00 (resp. $1$). The former case (2) is for $i_{v,S^{2}}=o_{v,S^{2}}=1$ and $i_{v,S^{1}\times S^{1}}=o_{v,S^{1}\times S^{1}}=0$ and we do not consider this case here.
@@
  • First we can see that for our resulting SSTF Morse function $f_{0}$ there exists exactly $c$ edges of $W_{f_{0}}$ such that the preimage ${q_{f_{0}}}^{-1}(t)$ of $t$ in the interior is diffeomorphic to the torus $S^{1}\times S^{1}$. Such edges are not mutually adjacent by our construction. The 1st Betti number of $W_{f_{0}}$ is greater than or equal to $a+b$. We can make the 1st Betti number an arbitrary integer $r\geq a+b$ by our construction. We remove the preimage of a small connected regular neighborhood of each of these $c$ edges. We have a $3$-dimensional manifold whose boundary is diffeomorphic to a disjoint union of $2c$ copies of $S^{2}$. We attach copies of the cylinder $S^{2}\times D^{1}$ instead to the boundary. This naturally yields another SSTF Morse function such that preimages of single points containing no singular points of it are disjoint unions of spheres. The resulting $3$-dimensional closed and connected manifold is diffeomorphic to the sphere $S^{3}$ in the case $a+b=0$ or a manifold diffeomorphic to a connected sum $\sharp(S^{1}\times S^{2})$ of $r\geq a+b$ copies of $S^{1}\times S^{2}$. This is regarded as a higher dimensional version of Theorem 1. We can understand the topology of the manifold $M$ from classical $3$-dimensional manifold theory for example. We may also have various ways of understanding. For example, on fundamental groups, we can also apply isomorphisms of fundamental groups of the manifold and the Reeb graph, discussed in [22, Corollary 4.8], [7], and [8, Corollary 4].
@@
We can deform such a Morse function to an SSTF Morse function $f_{0}:M\rightarrow\mathbb{R}$ by a suitable homotopy. We can also have such a function $f_{0}$ in such a way that to each local SSTF Morse function $f_{0,v}:X_{v}\rightarrow\mathbb{R}$ around each vertex $v$, the following handle attachment can be corresponded. Let $f_{0,v}(X_{v})=[t_{v,1},t_{v,2}]$ as in STEP 2.
@@
This is thanks to the topology of the manifold and fundamental arguments on $3$-dimensional manifold theory. The $3$-dimensional manifolds of the preimages of the local graphs in FIGURE 5 are all diffeomorphic to $(S^{1}\times D^{2})-{\rm Int}\ (D^{2}\sqcup D^{2})$. Here ${\rm Int}\ X$ means the interior $X-\partial X$ of the topological manifold $X$, the two disjoint disks $D^{2}\sqcup D^{2}$ are smoothly embedded in the interior of $S^{1}\times D^{2}$ and here its interior is removed. Note that similar local changes are presented in [16, Figure 4] and also used in [13] and that we cannot show this argument on the $3$-dimensional manifold from them directly. On the other hand, we can realize the local movements by applying suitable operations in [16, Figure 5] twice. This gives another proof for the local movements.
@@
Let a connected digraph $K_{g}$ be defined from a piecewise smooth function $g:K\rightarrow\mathbb{R}$ with the condition that the degree of a vertex where the function $g$ has a local extremum is of degree at most $2$. Then we have the following.
@@
    The function $f_{\rm m}:M\rightarrow\mathbb{R}$ is simple.
@@
The author would like to thank members of the research group of Osamu Saeki. The author would also like to thank people organizing and supporting Saga Souhatsu Mathematical Seminar (http://inasa.ms.saga-u.ac.jp/Japanese/saga-souhatsu.html): the author would also like to thank Inasa Nakamura for inviting the author as a speaker and letting him present [9, 11]. Discussing several Morse functions and topological properties of the manifolds of their domains with these persons and groups has motivated the author to challenge the present study.
@@
The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/) and this study is closely related to our study. Our study thanks them for the supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. Although he is not employed there, our study also thanks them.
Some of works by other researchers and this version may overlap in some of the contents due to the nature that our problems are natural in theory of Morse functions and applications to differential topology and that related mathematical studies are very fundamental and classical in some senses, for example. However the present version of our paper is presented independent of these work.